# LLM-based NLG Evaluation: Current Status and Challenges

### v2

@@ Abstract
Evaluating natural language generation (NLG) is a vital but challenging problem in artificial intelligence. Traditional evaluation metrics mainly capturing content (e.g. n-gram) overlap between system outputs and references are far from satisfactory, and large language models (LLMs) such as ChatGPT have demonstrated great potential in NLG evaluation in recent years. Various automatic evaluation methods based on LLMs have been proposed, including metrics derived from LLMs, prompting LLMs, and fine-tuning LLMs with labeled evaluation data. In this survey, we first give a taxonomy of LLM-based NLG evaluation methods, and discuss their pros and cons, respectively. We also discuss human-LLM collaboration for NLG evaluation. Lastly, we discuss several open problems in this area and point out future research directions.
**footnotetext: Equal contribution.
@@ 1 Introduction
Evaluating natural language generation (NLG) is a key but challenging issue. Traditional evaluation metrics like BLEU Papineni et al. (2002) and ROUGE Lin (2004) rely on the n-gram overlap between model outputs and references to measure its quality. They have been criticized for low correlation with human judgments Sulem et al. (2018), as surface-level matching cannot reliably evaluate text. After the rise of deep learning, model-based evaluation metrics like BERTScore Zhang et al. (2020) and BARTScore Yuan et al. (2021) have been continuously proposed and gradually adopted to evaluate the overall quality or various specific aspects of generated outputs (e.g., fluency, coherence, coverage, faithfulness, etc.). Although better than traditional metrics, their performance is still not satisfactory, and their application scope is very limited. For example, BERTScore is reference-based and cannot be used without a reference. With the emergence of large language models (LLMs) like InstructGPT Ouyang et al. (2022), they have achieved unprecedented effectiveness in following instructions, understanding content, and generating text. This inspired researchers to use LLMs for NLG evaluation. Although this is a research direction that only emerged in 2023, the past year has seen an enormous amount of research work. It is no exaggeration to say that NLG evaluation has been revolutionized by LLMs. This article will review the existing literature and provide suggestions for future research in this field.
Figure 1: Schematic representation of the four categories of LLM-based NLG evaluation.
This article mainly focuses on research that uses language models with over one billion parameters for NLG evaluation, with necessary references to earlier model-based evaluation metrics like BERTScore. To maintain focus, other types of generation like code generation and tasks involving images are not included in the scope of this article. As shown in Figure 1, according to how we utilize LLMs for NLG evaluation, we categorize the research work into four types:
  • LLM-derived Metrics2): developing/deriving evaluation metrics from embeddings or generation probabilities of LLMs.
  • Prompting LLMs3): directly inquiring existing LLMs via designed prompts which involve different elements for evaluation.
@@
  • Human-LLM Collaborative Evaluation5): leveraging distinctive strengths of both human evaluators and LLMs to achieve robust and nuanced evaluations in challenging domains through human-LLM collaboration.
We will review each type of evaluation methods and discuss the pros and cons respectively. Lastly, we will discuss future directions in this area (§6).
@@ 2 LLM-derived Metrics
2.1 Overview
Early model-based NLG evaluation methods and metrics, such as BERTScore and BARTScore, were primarily motivated by the capability of traditional pre-trained language models to generate high-quality texts. Recently, the advent of LLMs has prompted some research to adapt these ideas to stronger LLMs. Their inherent, powerful linguistic abilities are derived for direct NLG evaluation, which is expected to result in better performance. Such works can be categorized into two main types: embedding-based metrics and probability-based metrics.
@@ 2.2 Embedding-based Metrics
The embedding-based methods, like BERTScore, generally utilize representations of language models and thus compute the semantic similarity between the reference and the target text to evaluate, with different possible ways of implementation. Recent work ES et al. (2023) similarly obtains embeddings for the target text using the text-embedding-ada-002 model through the OpenAI API. The higher the corresponding similarity, the closer the target text aligns with the relevant requirements, indicating a higher quality.
@@ 2.3 Probability-based Metrics
To better utilize the knowledge inherent in language models, probability-based methods like BARTScore formulate text generation evaluation as conditional probability comparison, positing that the better the quality of the target text, the higher the likelihood that models should be able to generate it. Recently, GPTScore Fu et al. (2023a) has established tailored evaluation templates for each aspect to effectively guide multiple LLMs for NLG evaluation, including GPT3 Brown et al. (2020), OPT Zhang et al. (2022), and FLAN Chung et al. (2022). The generation probability is calculated under the condition of source input designed with customized prompts and evaluation aspects, which endows the GPTScore with better flexibility in evaluation. And similar methods have also been applied to the hallucination detection of the LLM-generated text Varshney et al. (2023) with three different attempts for calculating the probability score.
On the other hand, some works leverage the variation in probabilities under changed conditions as the evaluation metric. FFLM Jia et al. (2023) proposes to evaluate the faithfulness of the target text by calculating a combination of probability changes based on the intuition that the generation probability of a given text segment increases when more consistent information is provided, and vice versa. Similarly, DELTASCORE Xie et al. (2023) measures the quality of different story aspects according to the likelihood difference between pre- and post-perturbation states with LLMs including GPT-3.5 (text-davinci-003) that provide logits. They believe that the sensitivity to specific perturbations indicates the quality of related aspects, and their experiments demonstrate the effectiveness of their approach.
@@ 2.4 Pros and Cons
Efficiency. Compared to traditional metrics, LLM-derived evaluation methods are more time-consuming and require more computational resources, especially when adopting LLMs with quite large parameter scales. To address this, Eddine et al. (2022) proposes an approach to learning a lightweight version of LLM-derived metrics, and some fast LLM inference and serving tools like vLLM Kwon et al. (2023) have been launched. However, closed-source LLMs often do not make their parameters, representations, or logits public and available, thus making it impossible to apply LLM-derived methods to them.
Fairness. Sun et al. (2022) assesses the social bias across various metrics for NLG evaluation on six sensitive attributes: race, gender, religion, physical appearance, age, and socioeconomic status. Their findings reveal that model-based metrics carry noticeably more social bias than traditional metrics. Relevant biases can be categorized into two types: intrinsic bias encoded within pre-trained language models and extrinsic bias injected during the computation of similarity. Therefore, current LLM-derived methods may have similar issues.
@@ 3 Prompting LLMs
3.1 Overview
LLMs have demonstrated unprecedented instruction understanding and text generation abilities, which broadens researchers’ imagination of automatic evaluation for NLG. For a long time, human evaluation has been viewed as the gold standard for NLG evaluation. Recently, some studies claim that LLMs are on par with crowdsourcing annotators in several tasks Törnberg (2023); Gilardi et al. (2023); Ostyakova et al. (2023); Cegin et al. (2023). So, can LLMs simulate or even be an alternative to humans in the human evaluation of NLG? Or, do the practices in human evaluation or other evaluative tasks (e.g. competition race, paper review, etc.) inspire us to better NLG evaluation with LLMs? A large body of research and attempts based on prompting LLMs are guided by these ideas. In these studies, instructions and the text to be evaluated are completely expressed in the prompt given to LLMs, and the evaluation results are generated by them.
Human evaluation typically includes the following elements:
  • Evaluation Methods: The way the preferences of annotators are obtained, such as scoring and comparison.
  • Task Instructions: How the annotators should read or manipulate different parts to complete the annotation.
@@ 3.1 Overview
  • Role and interaction: The roles annotators play in the evaluation and the interactions between them.
The focus of the existing research can always be mapped analogously to one or more of these elements, and we organize them according to the elements they address. An example of prompting LLMs is shown in Figure 2.
Figure 2: An example of prompting LLMs to evaluate the aspect of consistency of the summary. There are task instructions, evaluation criteria, input content, and evaluation methods in the prompt, as well as the evaluation results, including the rating and explanation generated by LLMs.
@@ 3.2 Evaluation Methods
Scoring. Scoring is the most commonly used evaluation method in human evaluation for NLG, which is naturally applied to NLG evaluation by prompting LLMs. Chiang and Lee (2023a) has conducted relevant studies early, using a Likert scale from 1 to 5 to evaluate story generation and adversarial attacks with InstructGPT Ouyang et al. (2022) and ChatGPT  [^1] https://openai.com/blog/chatgpt/, showing that the evaluation results of LLMs are consistent with expert human evaluators. Kocmi and Federmann (2023b) discover GPT-3.5 and GPT-4 achieve the state-of-the-art accuracy of evaluating translation quality compared to human labels, outperforming all the results from the metric shard task of WMT22 Freitag et al. (2022). Wang et al. (2023a) experiment on five datasets across summarization, story generation, and data-to-text, and ChatGPT evaluators with a rating scale from 1 to 5 or 1 to 100 have the state-of-the-art or comparative correlations with human judgments in most settings, compared with prior metrics. Similar conclusions are also observed in open-domain dialogue response generation Lin and Chen (2023). Besides English, Mendonça et al. (2023) show that ChatGPT with simple rating prompts is a strong evaluator for multilingual dialogue evaluation, surpassing prior metrics based on encoders.
@@
Comparison. Different from absolute scoring, comparison refers to choosing the better of the two. Luo et al. (2023); Gao et al. (2023) use ChatGPT to compare the factual consistency of two summaries. AuPEL Wang et al. (2023d) evaluate personalized text generation from three aspects in the form of comparison with the PaLM 2 family Anil et al. (2023). According to Liusie et al. (2023), pairwise comparison is better than scoring when medium-sized LLMs (e.g. FlanT5 Chung et al. (2022) and Llama2 Touvron et al. (2023)) are adopted as evaluators.
@@
Ranking. Ranking can be viewed as an extended form of comparison. In comparison, only two examples are involved at a time, whereas in ranking, the order of more than two examples needs to be decided at once. Ji et al. (2023) use ChatGPT to rank five model-generated responses across several use cases at once, indicating the ranking preferences of ChatGPT align with those of humans to some degree. Similarly, GPTRank is a method to rank summaries in a list-wise manner Liu et al. (2023c). Moreover, Liu et al. (2023b) compare different evaluation methods in LLM-based summarization including scoring, comparison, and ranking, showing that the optimal evaluation method for each backbone LLM may vary.
@@
Boolean QA. Boolean QA requires LLMs to answer ”Yes” or ”No” to a question. It is adopted more in scenarios where human annotations are binary, such as grammaticality Hu et al. (2023), faithfulness of summaries and statements Luo et al. (2023); Gao et al. (2023); ES et al. (2023); Hu et al. (2023), factuality of generated text Fu et al. (2023b); Guan et al. (2023); Manakul et al. (2023), and answerability of generated questions Wang et al. (2023f).
@@
Error Analysis. Error Analysis refers to the evaluation of a text by looking for errors that occur in the text according to a set of predefined error categories. Multidimensional Quality Metrics (MQM) Jain et al. (2023) is an error analysis framework prevalent in machine translation evaluation. According to MQM, Lu et al. (2023); Kocmi and Federmann (2023a) use ChatGPT or GPT-4 to automatically detect translation quality error spans. BOOOOKSCORE Chang et al. (2023), an LLM-based evaluation metric, assesses the coherence of book summaries by identifying eight types of errors.

### v3

@@ Abstract
Evaluating natural language generation (NLG) is a vital but challenging problem in natural language processing. Traditional evaluation metrics mainly capturing content (e.g. n-gram) overlap between system outputs and references are far from satisfactory, and large language models (LLMs) such as ChatGPT have demonstrated great potential in NLG evaluation in recent years. Various automatic evaluation methods based on LLMs have been proposed, including metrics derived from LLMs, prompting LLMs, fine-tuning LLMs, and human-LLM collaborative evaluation. In this survey, we first give a taxonomy of LLM-based NLG evaluation methods, and discuss their pros and cons, respectively. Lastly, we discuss several open problems in this area and point out future research directions.
issue: 1**footnotetext: Equal contribution.
Contents
@@ 1 Introduction
The evaluation of natural language generation (NLG) is an important but challenging issue. The lack of a single standard answer and the presence of multiple quality criteria make evaluating NLG more challenging than other NLP tasks. For example, in news summarization, a good summary should capture the key information from the source document, remain faithful to the source document, and be expressed in logically coherent and fluent language, but there is not a single "correct" way to achieve this. The inherent difficulty of NLG evaluation means that human evaluation is always needed and regarded as the gold standard. However, due to the high cost and time-consuming nature of human evaluation, automatic evaluation metrics remain indispensable and play a crucial role in model development. Over the past two decades, many automatic evaluation metrics such as BLEU (Papineni et al., 2002) and BARTScore (Yuan, Neubig, and Liu, 2021) have been developed, but none have been fully satisfactory. Some studies (Sai et al., 2021; He et al., 2023a) have highlighted their deficiencies in robustness, such as insensitivities, biases, or even loopholes when evaluating challenging texts.
Recently, large language models (LLMs) have emerged and demonstrated unprecedented capacities in following instructions, understanding content, and generating text, which inspires researchers to utilize LLMs for NLG evaluation. Although this is a research direction that only emerged in 2023, the past year has seen an enormous amount of relevant work. While there have been surveys on automatic evaluation metrics or human evaluation practices in NLG evaluation (Celikyilmaz, Clark, and Gao, 2020; Hämäläinen and Al-Najjar, 2021; Zhou et al., 2022; Sai, Mohankumar, and Khapra, 2023; Gehrmann, Clark, and Sellam, 2023; Zhou, Ringeval, and Portet, 2023), none of them addresses LLM-based evaluation approach, and a comprehensive survey of this area is urgently needed.
The survey will mainly focus on research on LLM-based approaches for NLG evaluation, which involves language models with over one billion parameters. Moreover, it mainly considers the typical scope of NLG tasks where both input and output are natural languages including machine translation, text summarization, story generation, dialogue response generation, data-to-text, text simplification, paraphrase generation, grammatical error correction, and creative writing. Broader areas like evaluation of LLMs are not included (Zhuang et al., 2023; Ma et al., 2025) because this work focuses on LLMs used for evaluation, rather than the evaluation of LLMs’ capabilities. We search the literature on Google Scholar with an end date of June 2024 with keywords. Since this is a new direction that emerged in 2023, a considerable number of arXiv preprints are included in addition to papers published in *ACL venues or other related venues. About 100 pieces of work will be included. To maintain focus, this paper neither discusses datasets and benchmarks in NLG evaluation (Gehrmann et al., 2021; Kim et al., 2024b) nor analyzes evaluation metrics statistically (Ni’mah et al., 2023; Xiao et al., 2023).
Figure 1: Schematic representation of our proposed four categories of LLM-based NLG evaluation.
As shown in Figure 1, we categorize related studies into four categories according to how LLMs are utilized for NLG evaluation:
  • LLM-derived Metrics 2): developing or deriving evaluation metrics from embeddings or generation probabilities of LLMs.
  • Prompting LLMs 3): directly inquiring of existing LLMs via specific prompts and processes designed for evaluation.
@@
  • Human-LLM Collaborative Evaluation 5): leveraging distinctive strengths of both human evaluators and LLMs to achieve robust and nuanced evaluations through human-LLM collaboration.
LLMs have driven NLG evaluation toward a more human-centered direction, and the four categories we propose reflect this evolution: LLM-derived metrics are a continuation of traditional evaluation metrics and can only handle coarse-grained evaluation; prompting and fine-tuning methods enable users to express flexible evaluation requirements in natural language; collaborative evaluation takes it a step further, making it possible for humans and LLMs to leverage their strength respectively. We will review each type of evaluation method and discuss the pros and cons, respectively. Last but not least, we will provide our suggestions and conclusions, and discuss future directions in this area (§ 6).
It is worth stating that since LLM-based evaluation has shown unprecedented generality across NLG tasks, we do not summarize the literature for each task separately. Nevertheless, we will draw a list documenting all the approaches in this survey, indicating which NLG tasks each approach has been experimented on.
@@ 2 LLM-derived Metrics
LLM-derived metrics can be viewed as a continuation of early model-based NLG evaluation metrics such as BERTScore and BARTScore, replacing traditional pre-trained language models with stronger LLMs. Such works can be categorized into two main types: embedding-based metrics (ES et al., 2024) and probability-based metrics. The latter can be further divided into two categories based on different ways of using probabilities: directly converting the probabilities into scores (Fu et al., 2024; Varshney et al., 2023) and leveraging the variation in probabilities under changed conditions (Jia et al., 2023; Xie et al., 2023).
@@ 2.1 Embedding-based Metrics
[…]
@@ 2.2 Probability-based Metrics
To better utilize the knowledge inherent in language models, probability-based methods like BARTScore formulate text generation evaluation as conditional probability comparison, positing that the better the quality of the target text, the higher the likelihood that models should be able to generate it. Recently, GPTScore (Fu et al., 2024) has established tailored evaluation templates for each aspect to effectively guide multiple LLMs for NLG evaluation, including GPT3 (Wang et al., 2022), OPT (Zhang et al., 2022), and FLAN (Chung et al., 2024). The core idea of GPTScore is that a good generative language model is more likely to assign higher probabilities to high-quality text generated in response to a given instruction and context. Specifically, given a generative large language model $\theta$, context information $X$ (such as a source document), output text $Y=\{y_{1},y_{2},\dots,y_{m}\}$ containing $m$ tokens to be evaluated, and instruction $I$ that specifies the requirement for the LLMs to generate text that can flexibly correspond to different evaluation aspects (e.g., generating a factually consistent summary for the aspect of consistency), GPTScore is defined as:
[…]
Similarly, Murugadoss et al. (2025) score the task output $Y$ to be evaluated by its perplexity under the corresponding large language model $\theta$, given only the task context $X$. They believe this approach is unbiased by prompts, which transparently measures alignment with model training data. Furthermore, such methods have also been applied to the hallucination detection of the LLM-generated text (Varshney et al., 2023) with three different attempts for calculating the probability score.
On the other hand, some works leverage the variation in probabilities under changed conditions as the evaluation metric. FFLM (Jia et al., 2023) proposes to evaluate the faithfulness of the target text by calculating a combination of probability changes based on the intuition that the generation probability of a given text segment increases when more consistent information is provided, and vice versa. Similarly, DELTASCORE (Xie et al., 2023) measures the quality of different story aspects according to the likelihood difference between pre- and post-perturbation states with LLMs including GPT-3.5 (text-davinci-003) that provide logits. They believe that the sensitivity to specific perturbations indicates the quality of related aspects, and their experiments demonstrate the effectiveness of their approach.
@@ 2.3 Pros and Cons
Efficiency. Compared to traditional metrics, LLM-derived evaluation methods are more time-consuming and require more computational resources, especially when adopting LLMs with quite large parameter scales. To address this, Eddine et al. (2022) propose an approach to learning a lightweight version of LLM-derived metrics, and some fast LLM inference and serving tools like popular vLLM (Kwon et al., 2023) have been launched. vLLM improves memory utilization during inference through the PagedAttention algorithm, as well as the optimized memory management and batching strategies, thereby increasing LLMs’ generation throughput. However, closed-source LLMs often do not make their parameters, representations, or logits public and available, thus making it impossible to apply LLM-derived methods to them.
Fairness. Sun et al. (2022) assess the social bias across various metrics for NLG evaluation on six sensitive attributes: race, gender, religion, physical appearance, age, and socioeconomic status. Their findings reveal that model-based metrics carry noticeably more social bias than traditional metrics. Relevant biases can be categorized into two types: intrinsic bias encoded within pre-trained language models and extrinsic bias injected during the computation of similarity. Therefore, current LLM-derived methods may have similar issues.
@@ 3 Prompting LLMs
Related Work Evaluation Method NLG Task
Chiang and Lee (2023a) Scoring Story Generation
Wang et al. (2023a) Scoring Summarization, Data-to-text & Story Generation
Kocmi and Federmann (2023b) Scoring Translation
Lin and Chen (2023) Scoring Dialogue
Mendonça et al. (2023) Scoring Dialogue
Naismith, Mulcaire, and Burstein (2023) Scoring Discourse Generation
Liusie, Manakul, and Gales (2024) Scoring & Comparison Summarization, Dialogue & Data-to-text
Wang et al. (2023c) Comparison Personalized Text Generation
Ji et al. (2023) Ranking Open-end Text Generation
Liu et al. (2024b) Scoring, Ranking & Comparison Summarization
Wang, Funakoshi, and Okumura (2023) Boolean QA Question Generation
Manakul, Liusie, and Gales (2023) Boolean QA Fact Verification
Guan et al. (2024) Boolean QA Fact Verification
ES et al. (2024) Boolean QA Retrieval Augmented Generation
Kocmi and Federmann (2023a) Error Analysis Translation
Lu et al. (2023) Error Analysis Translation
Chang et al. (2024) Error Analysis Summarization
Table 1: Representative studies on prompting LLMs for NLG evaluation.
The remarkable generation abilities of LLMs have expanded the possibilities for NLG evaluation. For a long time, human evaluation has been viewed as the gold standard for NLG evaluation. Recently, some studies claim that LLMs are on par with crowdsourcing annotators in several tasks (Törnberg, 2023; Gilardi, Alizadeh, and Kubli, 2023; Ostyakova et al., 2023; Cegin, Simko, and Brusilovsky, 2023). This raises questions about whether LLMs could replace human evaluators. Studies in this area often involve feeding LLMs with detailed prompts that include both instructions and the text to be evaluated, with LLMs producing the evaluation outcomes. An example of prompting LLMs is shown in Figure 2. From this example, we can see that such a prompt is quite similar to the guidelines given to human evaluators. The main differences between this prompting method for LLMs and LLM-derived metrics are twofold: (1) LLM-derived metrics generally do not involve highly human-like prompts that require the LLM to perform an evaluation. (2) The evaluation results from prompting LLMs are typically generated directly by the LLM, whereas LLM-derived metrics require further transformation from embeddings and probabilities. We will describe existing works according to the five elements that they mainly focus on:
  • Evaluation Methods: The way the evaluation results of LLM evaluators are obtained, such as scoring and comparison.
  • Task Instructions: How LLM evaluators should read or manipulate different parts to complete the annotation.
@@
  • Role and interaction: The roles LLM evaluators play in the evaluation and the interactions between them.
Figure 2: An example of prompting LLMs to evaluate the aspect of consistency of the summary. There are role and interaction, task instructions, evaluation criteria, input content, and evaluation methods in the prompt, as well as the evaluation results, including the rating and explanation generated by LLMs.
@@ 3.1 Evaluation Methods
Diverse evaluation methods have been employed in prompting LLMs to obtain its preferences for the text to be evaluated: scoring, comparison, ranking, boolean QA, and error analysis (Table 1).
@@
Scoring. Scoring is the most commonly used evaluation method in human evaluation for NLG (van der Lee et al., 2021), and it is naturally applied to LLM-based evaluation. Chiang and Lee (2023a) have conducted relevant studies early, using a Likert scale from 1 to 5 to evaluate story generation and adversarial attacks with InstructGPT (Lee, 2024) and ChatGPT  [^1] https://openai.com/blog/chatgpt/, showing that the evaluation results of LLMs are consistent with expert human evaluators. Kocmi and Federmann (2023b) discover GPT-3.5 and GPT-4 achieve the state-of-the-art accuracy of evaluating translation quality compared to human labels with a rating scale from 1 to 5 or 0 to 100, outperforming all the results from the metric shard task of WMT22 (Freitag et al., 2022). Furthermore, Wang et al. (2023a) experiment on five datasets across summarization, story generation, and data-to-text, and ChatGPT with similar rating scales achieves the state-of-the-art or comparative correlations with human judgments in most settings, compared with prior metrics. Similar conclusions are also observed in open-domain dialogue response generation (Lin and Chen, 2023). Besides English, Mendonça et al. (2023) show that ChatGPT with simple rating prompts is a strong evaluator for multilingual dialogue evaluation, surpassing prior metrics based on encoders.
@@
Comparison. Different from absolute scoring, comparison refers to choosing the better of the two. Luo, Xie, and Ananiadou (2023); Gao et al. (2023) use ChatGPT to compare the factual consistency of two summaries. AuPEL (Wang et al., 2023c) evaluate personalized text generation from three aspects in the form of comparison with the PaLM 2 family (Anil et al., 2023). According to Liusie, Manakul, and Gales (2024), pairwise comparison is better than scoring when medium-sized LLMs (e.g. FlanT5 (Chung et al., 2024) and Llama2 (Touvron et al., 2023)) are adopted as evaluators.
@@
Ranking. Ranking can be viewed as an extended form of comparison. In comparison, only two examples are involved at a time, whereas in ranking, the order of more than two examples needs to be decided at once. Ji et al. (2023) use ChatGPT to rank five model-generated responses across several use cases at once, indicating the ranking preferences of ChatGPT align with those of humans to some degree. Similarly, GPTRank is a method to rank summaries in a list-wise manner (Liu et al., 2024b). Moreover, Liu et al. (2024a) compare different evaluation methods in LLM-based summarization including scoring, comparison, and ranking, showing that the optimal evaluation method for each backbone LLM may vary.
@@
Boolean QA. Boolean QA requires LLMs to answer "Yes" or "No" to a question. It is adopted more in scenarios where human annotations are binary, such as grammaticality (Hu et al., 2023), faithfulness of summaries and statements (Luo, Xie, and Ananiadou, 2023; Gao et al., 2023; ES et al., 2024; Hu et al., 2023), factuality of generated text (Fu et al., 2023; Guan et al., 2024; Manakul, Liusie, and Gales, 2023), and answerability of generated questions (Wang, Funakoshi, and Okumura, 2023).
@@
Error Analysis. Error Analysis refers to the evaluation of a text by looking for errors that occur in the text according to a set of predefined error categories. Multidimensional Quality Metrics (MQM) (Jain et al., 2023) is an error analysis framework prevalent in machine translation evaluation. According to MQM, Lu et al. (2023); Kocmi and Federmann (2023a) use ChatGPT or GPT-4 to automatically detect translation quality error spans. BOOOOKSCORE (Chang et al., 2024), an LLM-based evaluation metric, assesses the coherence of book summaries by identifying eight types of errors.